In a given normal distribution, the population mean is 80 and the population standard deviation is 2.5. Find the corresponding standard z – score of the following values.

1
Expert's answer
2022-05-23T16:50:14-0400

We can find z-score using this formula:

"Z=\frac{X-\mu}{\sigma}"

where X is observed value, "\mu" is mean of the sample, "\sigma" is the standard deviation.

We have: "\mu=80, \sigma=2.5"

So, z-value: "Z=\frac{X-80}{2.5}"

There is no X value in this problem, so you should just substitute X for your value and calculate z-score.

For example, if X=84, z-score is "Z=\frac {84-80}{2.5}=1.6"